Question 1 (50 points). You were recently assigned to manage a project for your company. You have constructed a network diagrShow all Calculations
Answer #1

Te = (To+4*Tp+Tm)/6

Variance = [( Tm-To)/6]2

Activity To Tm Tp Te Sigma Sigma^2
A 5 8 11 8 1 1
B 4 9 14 9 1.666667 2.777778
C 5 7 9 7 0.666667 0.444444
D 2 4 12 5 1.666667 2.777778
E 4 7 10 7 1 1

The critical path is ADE with expected project duration of 20 days

SD of critical path = ( 1+ 2.777+1)1/2 =2.185

Probability of completing the project in 21 days

z = x ( days in case) - mu ( avg completion time) / SD of critical path

z = 21-20 /2.185 = 0.4576

which corresponds to probability of 0.675

Probability of completing the project in 18 days

z = 18-20/2.185 = -0.9153

which corresponds to probability of 0.1814.